# From Citation Selection to Citation Absorption: A Measurement Framework for Generative Engine Optimization Across AI Search Platforms

### 28 Apr 2026 v1

@@ Abstract
Author and data note. Author order: Zhang Kai first and Yao Jingang second. Yao Jingang’s public profiles are GitHub https://github.com/yaojingang and X https://x.com/yaojingang. The public project repository is https://github.com/yaojingang/geo-citation-lab.
Counter-Intuitive Findings Placed Before the Main Argument
@@
Scientific status. This v2.0 manuscript is a formal pre-submission draft. It deliberately avoids fabricated p-values, confidence intervals, or regression coefficients. All numerical claims in the results section are descriptive and traceable to the public geo-citation-lab report and repository. Confirmatory inference specifications are included so that the final arXiv version can be populated after a fresh raw-CSV rerun under an explicit analysis script.
Introduction
@@
This paper is framed as an empirical measurement study. It does not claim access to hidden model internals, and it does not claim that observed content features causally force a generative engine to cite or use a page. It establishes a descriptive framework, documents reproducible patterns, and specifies the inferential models required for a final causal or quasi-causal extension.
Contributions
@@
    A scientific self-audit and reproducibility checklist designed to make the next arXiv submission defensible under independent review.
Related Work
Generative Engine Optimization
@@
Recent work has extended GEO toward AI search benchmarking, document-centric optimization, and diagnosis of citation failures [2,3]. The 2026 AgentGEO line of research emphasizes failure modes across retrieval, fetching, parsing, and generation [3]. This supports the premise that citation outcomes should be studied as a pipeline rather than as a single ranking event.
Generative Search, RAG, and Source Attribution
Generative search inherits elements from retrieval-augmented generation: retrieve candidate material, ground a response, and provide attribution [10,11]. Deployed answer engines introduce additional uncertainty because source selection, answer generation, and citation rendering are often black-box platform decisions. Studies of answer engines and source-cited responses document limitations such as hallucination, inaccurate citation, and mismatch between evidence and generated claims [4,7,8,9]. Citation-evaluation papers further show that citation quality needs sentence-level or claim-level analysis, since a source link can be present while failing to support the statement it is attached to [6].
@@
Why GEO Needs a New Dependent Variable
@@
Dataset and Provenance
Project Scope
The geo-citation-lab repository describes itself as a dataset and analysis pipeline for studying how AI search engines select and use citations [15]. The repository README identifies the original research author as Zhang Kai and the secondary report/open-source organization role as Yao Jingang [15]. This manuscript follows the user-specified author order: Zhang Kai first and Yao Jingang second.
The public snapshot includes 602 prompts, three platforms, 21,181 cleaned search-layer rows, 21,143 valid search-layer citations, 23,745 citation-level feature rows, 72 feature dimensions, and 18,151 successfully fetched citation pages. The reported overall fetch success rate is 76.44% [15].
@@
Cleaning and Caveats
The public report documents several cleaning caveats: repeated header rows in the ChatGPT CSV, normalization issues for ChatGPT A_news and A_technology naming, 15 missing ChatGPT prompt outputs after cleaning, noisy website-type values, and unknown values in country and language fields [15]. These details matter because cross-platform comparisons can be distorted if denominators are silently changed.
@@
Measurement Framework
Figure 1: The GEO measurement problem decomposes into citation selection and citation absorption. Source: author synthesis from the geo-citation-lab pipeline design.
Citation Selection
@@
Scientific Interpretation of the Influence Score
@@
Estimands and Model Specifications for the Final CSV Rerun
This manuscript reports descriptive estimands and specifies confirmatory models for the final version. The confirmatory models should be populated only after re-running the raw CSVs with a locked script, exporting model diagnostics, and documenting missingness.
@@
Identification Map: What Can and Cannot Be Claimed
@@
Claim level Example Supported by this draft? Required upgrade for stronger claim
@@
Results: Search Triggering and Citation Breadth
Figure 2: Search triggering is near universal, but citation breadth differs sharply by platform. Source: geo-citation-lab public report.
@@
Prompt Style and Language Effects
@@
Realistic and Extreme Scenarios
@@
Results: Source Selection
Figure 3: Source-type composition indicates that official, news, and vertical sources form the default candidate pool. Source: geo-citation-lab public report.
@@
Results: Citation Absorption
Figure 4: ChatGPT has substantially higher mean citation influence among fetched pages. Source: geo-citation-lab public report.
@@
Page Structure and Length
Figure 5: Ratio of top-quartile to bottom-quartile page attributes by influence score. Source: geo-citation-lab public report.
@@
Figure 6: Mean and median influence by word-count bin. Source: geo-citation-lab public report.
@@
Semantic Alignment and Evidence Genres
Figure 7: Reported correlations between independent features and influence_score. Variables used directly in the score definition are intentionally omitted from this figure. Source: geo-citation-lab public report.
@@
Figure 8: Mean influence uplift associated with evidence genres. Q&A format is the important negative case. Source: geo-citation-lab public report.
@@
Semantic Role and Usage Style
Figure 9: Semantic role determines how much a citation matters. Source: geo-citation-lab public report.
@@
Domain Type and Platform-Specific Correlations
@@
Figure 10: Platform-specific correlations indicate that each engine weights semantic and structural features differently. Source: geo-citation-lab public report.
@@
Platform Strategy Profiles as Descriptive Archetypes
@@
These archetypes should not be converted into permanent claims about vendor strategies. They are empirical profiles of the available snapshot. The final version should report model version, data-collection window, prompt execution order, and UI mode where possible. If those details cannot be reconstructed, the limitation should remain visible in the paper rather than hidden in a footnote.
Why the Negative Q&A Result Matters
@@
A plausible explanation is that Q&A formatting often creates short, isolated responses. Such pages may answer narrow questions but fail to provide the evidence density needed for synthesis across a more complex user prompt. Another explanation is taxonomy noise: some FAQ pages are thin support pages, and some non-Q&A pages are detailed explainers, government documents, or encyclopedic resources. A final regression should control for word count, structure, quality, relevance, and domain type before drawing a sharper conclusion.
@@
Industry and Question Type
Figure 11: Mean influence by question type. Source: geo-citation-lab public report.
@@
Mechanistic Interpretation: The Evidence-Container Hypothesis
@@
Evidence-Container Design: Operational Criteria
@@
Practical GEO Metrics Derived From the Framework
@@
Practical Implication Under Scientific Constraints
@@
Why This Differs From Classical SEO Advice
@@
Threats to Validity
Internal Validity
@@
Fetch-ok selection is also important. Absorption analysis excludes failed fetches. If failed pages systematically differ from successful pages, absorption estimates are conditional on fetchability. A final version should model fetch_ok as an intermediate selection outcome.
@@
Statistical Validity
The current manuscript uses descriptive statistics from the public report. A final confirmatory version should compute uncertainty intervals, correct for multiple comparisons, cluster errors by prompt and domain, and conduct robustness checks under alternative influence-score weights. The absence of these inferential quantities is a deliberate scientific choice in this draft, not an omission hidden as certainty.
Robustness and Validation Plan
@@
A scientifically defensible arXiv version should attach a reproducible notebook or script that starts from the repository CSV files and exports every table and figure. In addition, the final PDF should include a provenance table indicating which claims are direct observations, which are descriptive aggregations, and which are interpretive hypotheses.
Pre-Registered Analysis Plan for the Next Version
Before the final arXiv submission, the analysis should be pre-registered at the level of code and model specification. The primary selection outcome should be prompt-platform citation count. The primary absorption outcome should be influence score restricted to fetch-ok pages. Secondary outcomes should include selection trigger, domain-type category, semantic role, usage style, and content-genre indicators.
@@
Reproducibility and Data Availability
@@
A fully reproducible final version should include a version-pinned analysis script that starts from the repository CSV files, regenerates all descriptive tables and figures, records model versions for embedding and LLM-scoring steps, and exports the confirmatory inference tables described in the methodology section. Failed fetches should remain in the selection analysis and be excluded only for absorption analysis, with denominators reported explicitly.
The project repository is https://github.com/yaojingang/geo-citation-lab. Yao Jingang’s public profile links are included in the author block for traceability.
Conclusion
@@
The paper’s main scientific contribution is a measurement scaffold that separates exposure from answer influence. The main practical implication is evidence-container design. The main methodological warning is that descriptive GEO findings should not be prematurely converted into universal causal rules. The next version should rerun the raw CSV pipeline, add uncertainty estimates, and test whether controlled content rewrites can causally improve both selection and absorption.
Claim-Level Self-Audit
Claim Evidence status Limitation Action for final version
@@
Proposed Experimental Extension
The next research step is a controlled intervention study. Select a stratified sample of pages across industries and generate matched rewrites that vary only one feature family at a time: structure, evidence density, semantic alignment, source transparency, and page length. Submit matched prompts to the same platforms at the same timestamp window and compare selection and absorption outcomes.
@@
Compact Data Dictionary for the Final Reproduction Script
@@
Additional Claim-Level Notes for Reviewers
@@
The strongest remaining weakness is causal identification. A reviewer could reasonably ask whether high-influence pages are selected because they are high-quality, because they belong to favored domains, because they are longer, because they have better semantic alignment, or because the platform’s retrieval component surfaced them earlier. The next version should answer this with multivariate controls and controlled rewrites. Until then, the manuscript should be read as a rigorous measurement paper and hypothesis generator.

### 29 Apr 2026 v2

@@ Abstract.
Generative Engine Optimization, AI search, information retrieval, citation behavior, answer synthesis, source attribution, retrieval-augmented generation, content structure
Author and data note. Author order: Zhang Kai first, He Xinyue second, and Yao Jingang third. Yao Jingang public homepage information: GitHub https://github.com/yaojingang; X https://x.com/yaojingang. The public project repository used as the empirical source is https://github.com/yaojingang/geo-citation-lab. The repository and public report are cited for source traceability.
1. Counter-Intuitive Findings Placed Before the Main Argument
@@
Evidence standard. This manuscript avoids fabricated p-values, confidence intervals, or regression coefficients. All numerical claims in the results section are descriptive and traceable to the public geo-citation-lab report and repository. Confirmatory inference specifications are reported as a transparent analysis plan for future raw-CSV reruns under an explicit analysis script.
2. Introduction
@@
This paper is framed as an empirical measurement study. It does not claim access to hidden model internals, and it does not claim that observed content features causally force a generative engine to cite or use a page. It establishes a descriptive framework, documents reproducible patterns, and specifies the inferential models required for a causal or quasi-causal extension.
2.1. Contributions
@@
    A scientific self-audit and reproducibility checklist designed to support independent review and replication.
3. Related Work
3.1. Generative Engine Optimization
@@
Recent work has extended GEO toward AI search benchmarking, document-centric optimization, and diagnosis of citation failures [2,3,4,5,6]. Citation-repair work emphasizes that failures can occur across retrieval, fetching, parsing, attribution, and generation [3]. Agentic and structural GEO work further supports the premise that citation outcomes should be studied as a pipeline rather than as a single ranking event [5,6].
3.2. Generative Search, RAG, and Source Attribution
Generative search inherits elements from retrieval-augmented generation: retrieve candidate material, ground a response, and provide attribution [12,13,14,15]. Deployed answer engines introduce additional uncertainty because source selection, answer generation, and citation rendering are often black-box platform decisions. Studies of answer engines and source-cited responses document limitations such as hallucination, inaccurate citation, and mismatch between evidence and generated claims [7,8,9]. Citation-evaluation papers further show that citation quality needs sentence-level or claim-level analysis, since a source link can be present while failing to support the statement it is attached to [9].
@@ 3.2. Generative Search, RAG, and Source Attribution
3.3. Why GEO Needs a New Dependent Variable
@@
4. Dataset and Provenance
4.1. Project Scope
The geo-citation-lab repository describes itself as a dataset and analysis pipeline for studying how AI search engines select and use citations [16]. The repository README identifies the original research author as Zhang Kai and the secondary report/open-source organization role as Yao Jingang [16]. This manuscript lists Zhang Kai as first author, He Xinyue as second author, and Yao Jingang as third author. Yao Jingang’s public GitHub and X pages are included in the author note and references [18,19]. Repository-maintainer attribution is retained in the data-source references.
The public snapshot includes 602 prompts, three platforms, 21,181 cleaned search-layer rows, 21,143 valid search-layer citations, 23,745 citation-level feature rows, 72 feature dimensions, and 18,151 successfully fetched citation pages. The reported overall fetch success rate is 76.44% [16,17].
@@ 4.2. Prompt Architecture
4.3. Cleaning and Caveats
The public repository and report document several cleaning caveats: repeated header rows in the ChatGPT CSV, normalization issues for ChatGPT A_news and A_technology naming, 15 missing ChatGPT prompt outputs after cleaning, noisy website-type values, and unknown values in country and language fields [16,17]. These details matter because cross-platform comparisons can be distorted if denominators are silently changed.
@@ 4.3. Cleaning and Caveats
5. Measurement Framework
Figure 1. The GEO measurement problem decomposes into citation selection and citation absorption. Source: author synthesis from the geo-citation-lab pipeline design.
5.1. Citation Selection
@@ 5.2. Citation Absorption
5.3. Scientific Interpretation of the Influence Score
@@
5.4. Estimands and Model Specifications for a Raw CSV Rerun
This manuscript reports descriptive estimands and specifies confirmatory models for an expanded microdata analysis. The confirmatory models should be populated only after re-running the raw CSVs with a locked script, exporting model diagnostics, and documenting missingness.
@@ 5.4. Estimands and Model Specifications for a Raw CSV Rerun
5.5. Identification Map: What Can and Cannot Be Claimed
@@
Claim level Example Supported by this manuscript? Recommended evidence for stronger claim
@@
6. Results: Search Triggering and Citation Breadth
Figure 2. Search triggering is near universal, but citation breadth differs sharply by platform. Source: geo-citation-lab public report.
@@ 6. Results: Search Triggering and Citation Breadth
6.1. Prompt Style and Language Effects
@@
6.2. Realistic and Extreme Scenarios
@@
7. Results: Source Selection
Figure 3. Source-type composition indicates that official, news, and vertical sources form the default candidate pool. Source: geo-citation-lab public report.
@@ 7. Results: Source Selection
8. Results: Citation Absorption
Figure 4. ChatGPT has substantially higher mean citation influence among fetched pages. Source: geo-citation-lab public report.
@@ 8. Results: Citation Absorption
8.1. Page Structure and Length
Figure 5. Ratio of top-quartile to bottom-quartile page attributes by influence score. Source: geo-citation-lab public report.
@@ 8.1. Page Structure and Length
Figure 6. Mean and median influence by word-count bin. Source: geo-citation-lab public report.
@@
8.2. Semantic Alignment and Evidence Genres
Figure 7. Reported correlations between independent features and influence_score. Variables used directly in the score definition are intentionally omitted from this figure. Source: geo-citation-lab public report.
@@ 8.2. Semantic Alignment and Evidence Genres
Figure 8. Mean influence uplift associated with evidence genres. Q&A format is the important negative case. Source: geo-citation-lab public report.
@@
8.3. Semantic Role and Usage Style
Figure 9. Semantic role determines how much a citation matters. Source: geo-citation-lab public report.
@@ 8.3. Semantic Role and Usage Style
8.4. Domain Type and Platform-Specific Correlations
@@
Figure 10. Platform-specific correlations indicate that each engine weights semantic and structural features differently. Source: geo-citation-lab public report.
@@
8.5. Platform Strategy Profiles as Descriptive Archetypes
@@
These archetypes should not be converted into permanent claims about vendor strategies. They are empirical profiles of the available snapshot. A longitudinal replication should report model version, data-collection window, prompt execution order, and UI mode where possible. If those details cannot be reconstructed, the limitation should remain visible in the paper rather than hidden in a footnote.
8.6. Why the Negative Q&A Result Matters
@@
A plausible explanation is that Q&A formatting often creates short, isolated responses. Such pages may answer narrow questions but fail to provide the evidence density needed for synthesis across a more complex user prompt. Another explanation is taxonomy noise: some FAQ pages are thin support pages, and some non-Q&A pages are detailed explainers, government documents, or encyclopedic resources. A follow-up regression should control for word count, structure, quality, relevance, and domain type before drawing a sharper conclusion.
@@
8.7. Industry and Question Type
Figure 11. Mean influence by question type. Source: geo-citation-lab public report.
@@ 8.7. Industry and Question Type
9. Mechanistic Interpretation: The Evidence-Container Hypothesis
@@
9.1. Evidence-Container Design: Operational Criteria
@@
9.2. Practical GEO Metrics Derived From the Framework
@@
9.3. Practical Implication Under Scientific Constraints
@@
9.4. Why This Differs From Classical SEO Advice
@@
10. Threats to Validity
10.1. Internal Validity
@@
Fetch-ok selection is also important. Absorption analysis excludes failed fetches. If failed pages systematically differ from successful pages, absorption estimates are conditional on fetchability. A follow-up analysis should model fetch_ok as an intermediate selection outcome.
@@ 10.2. External Validity
10.3. Statistical Validity
The current manuscript uses descriptive statistics from the public report. A confirmatory microdata analysis should compute uncertainty intervals, correct for multiple comparisons, cluster errors by prompt and domain, and conduct robustness checks under alternative influence-score weights. The absence of these inferential quantities is a deliberate scientific choice in the current manuscript, not an omission hidden as certainty.
11. Robustness and Validation Plan
@@
A fully reproducible research release should attach a notebook or script that starts from the repository CSV files and exports every table and figure. The manuscript should also include a provenance table indicating which claims are direct observations, which are descriptive aggregations, and which are interpretive hypotheses.
11.1. Confirmatory Analysis Plan
The confirmatory analysis should be specified at the level of code and model design before inferential claims are reported. The primary selection outcome should be prompt-platform citation count. The primary absorption outcome should be influence score restricted to fetch-ok pages. Secondary outcomes should include selection trigger, domain-type category, semantic role, usage style, and content-genre indicators.
@@ 12. Responsible GEO and Ethics
13. Reproducibility and Data Availability
@@
A fully reproducible research release should include a version-pinned analysis script that starts from the repository CSV files, regenerates all descriptive tables and figures, records model versions for embedding and LLM-scoring steps, and exports the confirmatory inference tables described in the methodology section. Failed fetches should remain in the selection analysis and be excluded only for absorption analysis, with denominators reported explicitly.
The project repository is https://github.com/yaojingang/geo-citation-lab. Repository and public-report links are retained for source traceability.
14. Conclusion
@@
The paper’s main scientific contribution is a measurement scaffold that separates exposure from answer influence. The main practical implication is evidence-container design. The main methodological warning is that descriptive GEO findings should not be prematurely converted into universal causal rules. Follow-up work should rerun the raw CSV pipeline, add uncertainty estimates, and test whether controlled content rewrites can causally improve both selection and absorption.
Appendix A Claim-Level Self-Audit
Claim Evidence status Limitation Recommended follow-up action
@@ Appendix A Claim-Level Self-Audit
Appendix B Proposed Experimental Extension
A natural research extension is a controlled intervention study. Select a stratified sample of pages across industries and generate matched rewrites that vary only one feature family at a time: structure, evidence density, semantic alignment, source transparency, and page length. Submit matched prompts to the same platforms at the same timestamp window and compare selection and absorption outcomes.
@@ Appendix B Proposed Experimental Extension
Appendix C Compact Data Dictionary for the Reproduction Script
@@
Appendix D Reproducibility and Release Checklist
This checklist summarizes the manuscript’s reproducibility posture. It is written as a release-oriented research checklist rather than as a platform-specific filing checklist.
Item Current handling Status
Author metadata Author order is Zhang Kai first, He Xinyue second, and Yao Jingang third; Yao public profile links and repository links are retained for traceability. Included
Public dataset attribution The geo-citation-lab repository and public report are cited as the data source. Included
Evidence boundaries Descriptive claims, interpretive hypotheses, and proposed confirmatory models are separated in the text. Included
Denominator discipline Search-layer and fetch-ok denominators are reported separately. Included
Circularity control Influence-score components are excluded from primary explanatory language. Included
Microdata inference Bootstrap intervals, clustered standard errors, and multivariate models are specified as a follow-up microdata analysis. Recommended extension
Temporal replication Repeating the prompt battery over time is specified as a longitudinal robustness extension. Recommended extension
Appendix E Additional Claim-Level Notes
@@
The strongest remaining weakness is causal identification. A reviewer could reasonably ask whether high-influence pages are selected because they are high-quality, because they belong to favored domains, because they are longer, because they have better semantic alignment, or because the platform’s retrieval component surfaced them earlier. Follow-up work should answer this with multivariate controls and controlled rewrites. Until then, the manuscript should be read as a rigorous measurement paper and hypothesis generator.
Appendix F References
[1] Aggarwal, P., Murahari, V., Rajpurohit, T., Kalyan, A., Narasimhan, K., and Deshpande, A. (2024). GEO: Generative Engine Optimization. arXiv:2311.09735. Accepted to KDD 2024.
[2] Chen, M., Wang, X., Chen, K., and Koudas, N. (2025). Generative Engine Optimization: How to Dominate AI Search. arXiv:2509.08919.
[3] Tian, Z., Chen, Y., Tang, Y., Liu, J., and Jia, R. (2026). Diagnosing and Repairing Citation Failures in Generative Engine Optimization. arXiv:2603.09296.
[4] Liu, Z., and Xu, P. (2026). Think Before Writing: Feature-Level Multi-Objective Optimization for Generative Citation Visibility. arXiv:2604.19113.
[5] Yuan, J., Wang, J., Wang, Z., Sun, Q., Wang, R., and Li, J. (2026). AgenticGEO: A Self-Evolving Agentic System for Generative Engine Optimization. arXiv:2603.20213.
[6] Yu, J., Yang, M., Ding, Y., and Sato, H. (2026). Structural Feature Engineering for Generative Engine Optimization: How Content Structure Shapes Citation Behavior. arXiv:2603.29979.
[7] Narayanan Venkit, P., Laban, P., Zhou, Y., Mao, Y., and Wu, C.-S. (2024). Search Engines in an AI Era: The False Promise of Factual and Verifiable Source-Cited Responses. arXiv:2410.22349.
[8] Yang, K.-C. (2025). News Source Citing Patterns in AI Search Systems. arXiv:2507.05301.
[9] Xu, Y., Qi, P., Chen, J., Liu, K., Han, R., Liu, L., Min, B., Castelli, V., Gupta, A., and Wang, Z. (2025). CiteEval: Principle-Driven Citation Evaluation for Source Attribution. arXiv:2506.01829.
[10] Qian, H., Fan, Y., Zhang, R., and Guo, J. (2024). On the Capacity of Citation Generation by Large Language Models. arXiv:2410.11217.
[11] Kirsten, E., Grosse Perdekamp, J., Upadhyay, M., Gummadi, K. P., and Zafar, M. B. (2025). Characterizing Web Search in The Age of Generative AI. arXiv:2510.11560.
[12] Lewis, P., Perez, E., Piktus, A., Petroni, F., Karpukhin, V., Goyal, N., et al. (2020). Retrieval-Augmented Generation for Knowledge-Intensive NLP Tasks. Advances in Neural Information Processing Systems.
[13] Karpukhin, V., Oguz, B., Min, S., Lewis, P., Wu, L., Edunov, S., Chen, D., and Yih, W.-t. (2020). Dense Passage Retrieval for Open-Domain Question Answering. EMNLP.
[14] Nakano, R., et al. (2021). WebGPT: Browser-assisted question-answering with human feedback. arXiv:2112.09332.
[15] Menick, J., et al. (2022). Teaching language models to support answers with verified quotes. arXiv:2203.11147.
[16] geo-citation-lab repository. (2026). A dataset and analysis pipeline for studying how AI search engines select and use citations. GitHub: https://github.com/yaojingang/geo-citation-lab. Accessed April 28, 2026.
[17] geo-citation-lab final report. (2026). Overseas GEO Research Long Report, recalculated version. https://yaojingang.github.io/geo-citation-lab/04-repet/final_report.html. Accessed April 28, 2026.
[18] Yao Jingang. (2026). GitHub profile. https://github.com/yaojingang. Accessed April 29, 2026.
[19] Yao Jingang. (2026). X profile. https://x.com/yaojingang. Accessed April 29, 2026.